OpenStudy (anonymous):

he average of 3 numbers is 27. If a 4th number is added the new average is 22. What number was added?

OpenStudy (anonymous):

the answer is 7

OpenStudy (anonymous):

Angles X and Y are congruent. and . Find the degree measure of angle X.

OpenStudy (anonymous):

thaT makes no sense is that all the info

OpenStudy (whpalmer4):

if the average of the 3 numbers is 27, we can pretend that the numbers were each 27 — 3 27s obviously average out to 27, right? then we have (27 + 27 + 27 + x) / 4 = 22 81+x = 88 x = 88-81 x=7